OpenStudy (anonymous):

Solve the quadratic equation 5x2 - 125 = 0.

OpenStudy (anonymous):

Add 125 to both sides and then divide by 5. As a last step, take the positive and negative square roots.

OpenStudy (anonymous):

try \(5x^2=125\) then divide by 5 and finally take the square root, don't forget the \(\pm\)

OpenStudy (anonymous):

5x^2 - 125 = 0 -> 5x^2 = 125 x^2 = 125/5 = 25 -> x = +- sqrt(25) x = +- 5

OpenStudy (anonymous):

All good now, @Stephensparadox ?

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

Good luck in all of your studies and thx for the recognition! @Stephensparadox